# Gluon Condensate via Dirac Spectral Density:

### v2

@@ Abstract
Quark and gluon scalar densities, $\langle\bar{\psi}\psi\rangle$ and $\langle F^{2}\rangle$, reflect the degree of scale-invariance violations in SU(N) gauge theories with fundamental quarks. It is known that $\langle\bar{\psi}\psi\rangle$ can be usefully scale-decomposed via spectral density $\rho(\lambda)$ of Dirac modes. Here I give such formula for $\langle F^{2}\rangle$, which reveals that gluon condensate is a strictly UV quantity. For the recently-found IR phase [1, 2], where the infrared (IR) degrees of freedom separate out and become independent of the system’s bulk, it implies that $\langle F^{2}\rangle$ due to this IR part vanishes. Its glue thus doesn’t contribute to scale anomaly of the entire system and is, in this sense, scale invariant consistently with the original claim. Associated formulas are used us to define IR decoupling of glue, which may serve as an alternative indicator of IR phase transition. Using the simplest form of coherent lattice QCD, we express the effective action of full QCD entirely via Dirac spectral density.
@@
with $\lambda\!=\!\lambda_{\scriptscriptstyle{\mathrm{R}}}\!+\!i\lambda_{\scriptscriptstyle{\mathrm{I}}}$ a complex eigenvalue of lattice Dirac operator $D$ defining quark dynamics in a regularized theory (UV cutoff $1/a$; IR cutoff $1/L$). This $D$ is chosen such that the associated dimensionless constant $c_{\scriptscriptstyle{\mathrm{S}}}$ is non-zero (see below). Quantity $\rho_{s}(\lambda)\!=\!\langle n(\lambda,d{\cal S})\rangle/(V_{4}\,d{\cal S})$ is the surface spectral density of $D$, with $n(\lambda,d{\cal S})$ the number of eigenvalues in the area $d{\cal S}\!=\!d\lambda_{\scriptscriptstyle{\mathrm{R}}}d\lambda_{\scriptscriptstyle{\mathrm{I}}}$ around $\lambda$, and $V_{4}=L^{3}/T$. [^2] Notation ${\mathbb{R}}^{2}[{\mathbb{C}}]$ conveys that components in the domain ${\mathbb{R}}^{2}$ have meaning in the complex plane $\mathbb{C}$}. Dimension of $\rho_{s}(\lambda)$ is $a^{-2}$ instead of $a^{-3}$ in the standard definition. Effective density $\rho_{s}^{\scriptscriptstyle{\text{ef}}}\equiv\rho_{s}\!-\!\rho_{s0}$ subtracts that of the free field. [^3] We will see that in the glue case the free-field subtraction is necessary to merely define the operator. In the quark case it removes the leading UV divergence at non-zero $m$, which keeps the two expressions symmetric in this sense. Note that $\rho_{s}^{\scriptscriptstyle{\text{ef}}}=\rho_{s}^{\scriptscriptstyle{\text{ef}}}(\lambda,a,L)$ and $\int_{{\mathbb{R}}^{2}[{\mathbb{C}}]}d{\cal S}\;\rho_{s}^{\scriptscriptstyle{\text{ef}}}(\lambda)=0$.
@@
To that end, it is important to realize that (i) and (ii) are in fact connected [2, 4]. For example, while the IR-bulk separation is very suggestive of IR-bulk decoupling, it is not sufficient. But (ii) offers a dynamical reason for it which is easy to see in thermal IR phase at $N_{f}\!=\!0$. Indeed, assume that, upon entering the IR phase, this theory of glue becomes scale invariant below some energy $\Lambda_{\scriptscriptstyle{\text{\rm IR}}}<T$. The gauge coupling then stops running at that scale, which requires non-analyticities in the internal structure of the theory. These non-analyticities can then also facilitate the IR-bulk decoupling. Conversely, non-analyticities generated by breaking-off the IR into a physically distinct component can make scale invariance in IR possible. [^6] Thermal onset of these internal non-analyticities is predicted to occur at $T\!=\!T_{\scriptscriptstyle{\text{\rm IR}}}$ of IR phase transition [4, 5] because they induce non-analytic $T$-dependence of physical observables at that point.
@@
Few remarks are important here. (i) It is implicitly understood that the above definition doesn’t depend on the choice of $t_{\text{ref}}$. (ii) Regarding $t_{\text{dc}}$, the available evidence [4, 5, 6, 7] embodied in the metal-to-critical scenario [5] favors the simplest possibility that $t_{\text{dc}}\!=\!t_{3}\!=\!t_{A}$ (see also Ref. [24]), where $t_{A}\!>\!0$ is the Anderson-like point of Refs [27, 28, 29]. However, studies directly confirming the above form of IR-Bulk decoupling, as well as direct computation of $t_{\text{dc}}$ are yet to be performed. (iii) It is consistent with the current knowledge that decoupling Eq. (16) may occur already before taking the continuum limit, i.e. even at non-zero $a$. (iv) Detection of IR phase (or even its definition) via the presence of the above IR-Bulk decoupling may be a fruitful approach to its numerical investigation. (v) Normalized Pearson correlations of ${\cal F}^{2}$ segments, such as those entering Eq. (16) can be expressed via via spectral values $\sigma_{i}(U)$ alone. The general expression is
@@
where $\lambda$ is a complex variable and the notation was introduced in connection with Eq. (2). Thus, the defining object of the theory, the action, is scale-decomposed and expressed in terms of Dirac spectral density. There is another hidden constant in the above expression, isolated by factoring $am_{q}$ in log terms. Discarding it puts the effective action into the form
@@
The outlook on $\rho(\lambda)$ as the distribution of DOFs over scales was pervasive in works that led to the discovery of IR phase [1, 2, 4, 5]. Definition of the phase is also expressed via IR asymptotics of $\rho(\lambda)$ [2]. But its key features, such as scale invariance and IR-Bulk separation, were always chiefly attributed to gluonic rather than quark degrees of freedom. This was not contradictory with the explanation that matrix elements of the Dirac operator are in fact gauge-covariant glue operators, which makes $\rho(\lambda)$ the unusual (scale-dependent) but glue operator as well. Here we made the needed direct link between Dirac spectral density and the aforementioned driving glue effects.
Our formulas have several applications, particularly in relation to IR phase where these considerations first arose. Here we discussed the question of glue scale invariance of the IR part and the precise definition of IR-Bulk decoupling. They are all based on our formulas featuring Dirac spectral density, which also possibly opens new ways of numerical evaluation. Our considerations are most potent in regularizations known as coherent lattice QCD [13, 14], whose full effective actions can also be expressed via Dirac spectral density.

### v3

@@ Abstract
Quark and gluon scalar densities, $\langle\bar{\psi}\psi\rangle$ and $\langle F^{2}\rangle$, reflect the degree of scale-invariance violations in SU(N) gauge theories with fundamental quarks. It is known that $\langle\bar{\psi}\psi\rangle$ can be usefully scale-decomposed via spectral density $\rho(\lambda)$ of Dirac modes. Here we give such formula for $\langle F^{2}\rangle$, which reveals that gluon condensate is a strictly UV quantity. For the recently-found IR phase [1, 2], where the infrared (IR) degrees of freedom separate out and become independent of the system’s bulk, it implies that $\langle F^{2}\rangle$ due to this IR part vanishes. Its glue thus doesn’t contribute to scale anomaly of the entire system and is, in this sense, scale invariant consistently with the original claim. Associated formulas are used to define IR decoupling of glue, which may serve as an alternative indicator of IR phase transition. Using the simplest form of coherent lattice QCD, we express the effective action of full QCD entirely via Dirac spectral density.
@@
with $\lambda\!=\!\lambda_{\scriptscriptstyle{\mathrm{R}}}\!+\!i\lambda_{\scriptscriptstyle{\mathrm{I}}}$ a complex eigenvalue of lattice Dirac operator $D$ defining quark dynamics in a regularized theory (UV cutoff $1/a$; IR cutoff $1/L$). This $D$ is chosen such that the associated dimensionless constant $c_{\scriptscriptstyle{\mathrm{S}}}$ is non-zero (see below). Quantity $\rho_{s}(\lambda)\!=\!\langle n(\lambda,d{\cal S})\rangle/(V_{4}\,d{\cal S})$ is the surface spectral density of $D$, with $n(\lambda,d{\cal S})$ the number of eigenvalues in the area $d{\cal S}\!=\!d\lambda_{\scriptscriptstyle{\mathrm{R}}}d\lambda_{\scriptscriptstyle{\mathrm{I}}}$ around $\lambda$, and $V_{4}=L^{3}/T$. [^2] Denoting the integration domain ${\mathbb{R}}^{2}[{\mathbb{C}}]$ conveys that components of ${\mathbb{R}}^{2}$ have meaning in the complex plane $\mathbb{C}$}. Dimension of $\rho_{s}(\lambda)$ is $a^{-2}$ instead of $a^{-3}$ in the standard definition. Effective density $\rho_{s}^{\scriptscriptstyle{\text{ef}}}\equiv\rho_{s}\!-\!\rho_{s0}$ subtracts that of the free field. [^3] We will see that in the glue case the free-field subtraction is necessary to merely define the operator. In the quark case it removes the leading UV divergence at non-zero $m$, which keeps the two expressions symmetric in this sense. Note that $\rho_{s}^{\scriptscriptstyle{\text{ef}}}=\rho_{s}^{\scriptscriptstyle{\text{ef}}}(\lambda,a,L)$ and $\int_{{\mathbb{R}}^{2}[{\mathbb{C}}]}d{\cal S}\;\rho_{s}^{\scriptscriptstyle{\text{ef}}}(\lambda)=0$.
@@
To that end, it is important to realize that (i) and (ii) are in fact connected [2, 4]. For example, while the IR-bulk separation is very suggestive of IR-bulk decoupling, it is not sufficient for the claim. But (ii) offers a dynamical reason for it which is easy to see in thermal IR phase at $N_{f}\!=\!0$. Indeed, assume that, upon entering the IR phase, this theory of glue becomes scale invariant below some energy $\Lambda_{\scriptscriptstyle{\text{\rm IR}}}<T$. The gauge coupling then stops running at that scale, which requires non-analyticities in the internal structure of the theory. These non-analyticities can then also facilitate the IR-bulk decoupling. Conversely, non-analyticities generated by breaking-off the IR into a physically distinct component can make scale invariance in IR possible. [^6] Thermal onset of these internal non-analyticities is predicted to occur at $T\!=\!T_{\scriptscriptstyle{\text{\rm IR}}}$ of IR phase transition [4, 5] because they induce non-analytic $T$-dependence of physical observables at that point.
@@
Few remarks are important here. (i) It is implicitly understood that the above definition doesn’t depend on the choice of $t_{\text{ref}}$. (ii) Regarding $t_{\text{dc}}$, the available evidence [4, 5, 6, 7] embodied in the metal-to-critical scenario [5] favors the simplest possibility that $t_{\text{dc}}\!=\!t_{3}\!=\!t_{A}$ (see also Ref. [24]), where $t_{A}\!>\!0$ is the Anderson-like point of Refs [27, 28, 29]. However, studies directly confirming the above form of IR-Bulk decoupling, as well as direct computation of $t_{\text{dc}}$ are yet to be performed. (iii) It is consistent with the current knowledge that decoupling Eq. (16) may occur already before taking the continuum limit, i.e. even at non-zero $a$. (iv) Detection of IR phase (or even its definition) via the presence of the above IR-Bulk decoupling may be a fruitful approach to its numerical investigation. (v) Normalized Pearson correlations of ${\cal F}^{2}$ segments, such as those entering Eq. (16) can be expressed via spectral values $\sigma_{i}(U)$ alone. The general expression is
@@
where $\lambda$ is a complex variable and the notation was introduced in connection with Eq. (2). Thus, the defining object of the theory, the action, is scale-decomposed and expressed in terms of Dirac spectral density. There is another hidden constant in the above expression, isolated by factoring $am_{q}$ in log terms. Discarding this constant leads to the effective action in the form
@@
The outlook on $\rho(\lambda)$ as the distribution of DOFs over scales was pervasive in works that led to the discovery of IR phase [1, 2, 4, 5]. Definition of the phase is also expressed via IR asymptotics of $\rho(\lambda)$ [2]. But its key features, such as scale invariance and IR-Bulk separation, were always chiefly attributed to gluonic rather than quark degrees of freedom. This was not contradictory with the explanation that matrix elements of the Dirac operator are in fact gauge-covariant glue operators, which makes $\rho(\lambda)$ the unusual (scale-dependent) glue operator as well. Here we made the needed direct link between Dirac spectral density and the aforementioned driving glue effects.
Our formulas have several applications, particularly in relation to IR phase where these considerations first arose. Here we discussed the question of glue scale invariance of the IR part and the precise definition of IR-Bulk decoupling. They are all based on our formulas featuring Dirac spectral density, which possibly also opens new ways of numerical evaluation. Our considerations are most potent in regularizations known as coherent lattice QCD [13, 14], whose full effective actions can also be expressed via Dirac spectral density.
Finally, we wish to emphasize that although the use of overlap Dirac operator for expressing gluon condensate (Eq. (7)) has many theoretical advantages, simpler operators may be preferable in numerical studies. In particular, the formula for all lattice Dirac operators with $c_{\scriptscriptstyle{\mathrm{S}}}\!\neq\!0$ is given in (2). For example, simple Wilson-like Dirac operators with $c_{\scriptscriptstyle{\mathrm{S}}}\!\neq\!0$ can be considered.